### v1

# On average population levels for models with directed diffusion in heterogeneous environments

Abstract
In 2006 (J. Differential Equ.), Lou proved that, once the intrinsic growth rate $r$ in the logistic model is proportional to the spatially heterogeneous carrying capacity $K$ ($r=K^{1}$), the total population under the regular diffusion exceeds the total of the carrying capacity. He also conjectured that the dependency of the total population on the diffusion coefficient is unimodal, increasing to its maximum and then decreasing to the asymptote which is the total of the carrying capacity. DeAngelis et al (J. Math. Biol. 2016) argued that the prevalence of the population over the carrying capacity is only observed when the growth rate and the carrying capacity are positively correlated, at least for slow dispersal. Guo et al (J. Math. Biol. 2020) justified that, once $r$ is constant ($r=K^{0}$), the total population is less than the cumulative carrying capacity. Our paper fills up the gap for when $r=K^{\lambda}$ for any real $\lambda$, disproving an assumption that there is a critical $\lambda^{\ast}\in(0,1)$ at which the tendency of the prevalence of the carrying capacity over the total population size changes, demonstrating instead that the relationship is more complicated. In addition, we explore the dependency of the total population size on the diffusion coefficient when the third parameter of the dispersal strategy $P$ is involved: the diffusion term is $d\Delta(u/P)$, not just $d\Delta u$, for any $\lambda$. We outline some differences from the random diffusion case, in particular, concerning the profile of the total population as a function of the diffusion coefficient.
AMS subject classification: 92D25, 35J25, 92D40, 35J60, 35B40
Keywords: spatial heterogeneity, random and directed diffusion, average population level, mathematical ecology
1 Introduction
The idea to include diffusion for a more comprehensive description of dispersal in population ecology was a significant breakthrough in analysing invasions and spatial species interactions [27]. Together with random walks approach, application of diffusion became a cornerstone in analysing populations movements and their effect [14, 15, 17, 24]. Since different species disperse in unique ways, their specific dispersal mechanisms must be accounted for when incorporating spatial heterogeneity into time-dependent mathematical models [5]. The effects of spatial heterogeneity were explored in [19, 20, 21]. The environment variability was strongly connected to the notion of the carrying capacity which was actively discussed [26, 28].
This paper studies the relationship between the carrying capacity of the environment and the total population size of a single species adopting different dispersal strategies. This connection was explored in [14, 15, 18, 25], where population dynamics was modelled using the logistic equation with random diffusion
$\begin{cases}\frac{\partial u}{\partial t}(x,t)=d\Delta u(t,x)+r(x)u(t,x)\left(1-\frac{u(t,x)}{K(x)}\right)=0,&x\in\Omega,~~t>0\\ u(x,0)=u_{0}(x),&x\in\Omega\\ \frac{\partial u}{\partial n}(t,x)=0,&x\in\partial\Omega,~~t>0.\end{cases}$ (1.1)
Here $r(x)$ and $K(x)$ are time-independent positive on $\bar{\Omega}$ functions representing the intrinsic growth rate and the carrying capacity, respectively.
The disadvantage of considering that species diffuse randomly as in (1.1) is that, as the diffusion becomes more important in our model (when $d$ becomes large enough), the solution of (1.2) goes away from the carrying capacity profile [8, 10, 11, 13]. This leads to the counterintuitive outcome that in the competition of several species identical in all the parameters but the value of the diffusion coefficient $d$ in (1.1), the slower diffuser wins [16]. Multiple studies in mathematical biology indicated that it is reasonable to consider that populations diffuse in an ideal free dispersal framework, characterized by assuming populations can diffuse as they wish, but they choose to disperse from places with lower suitability to places where their chances of survival increase [2, 6, 9]. Populations choosing an ideal free distribution are found to be harder to be invaded by competing species sharing their habitat [4, 8, 12, 22, 23], and therefore, a directed diffusion based on the heterogeneity of the environment becomes more present on the modelling of population dynamics, see [7, 10] for a detailed overview.
Let $u_{d}$ denote a stationary solution to (1.1), i.e., let $u_{d}$ be a solution to
$\begin{cases}\displaystyle d\Delta u(x)+r(x)u(x)\left(1-\frac{u(x)}{K(x)}\right)=0,&x\in\Omega,\\ \displaystyle\frac{\partial u}{\partial n}(x),&x\in\partial\Omega.\end{cases}$ (1.2)
Then, as the diffusion coefficient tends to zero, $u_{d}$ tends to the carrying capacity at every point of the domain $\Omega$, and as $d\to+\infty$, the stationary solution $u_{d}$ is distributed uniformly over $\Omega$ [14, 25]. However, a problem still under investigation is the relation between the total stationary population for $0<d<+\infty$ and the total carrying capacity.
Lou [25] demonstrated that for $r(x)=\alpha K(x)$ for some constant $\alpha>0$, the total population size of the solution $u_{d}$ to (1.2) exceeds the total carrying capacity for every $d\in(0,+\infty)$, i.e.,
$\int_{\Omega}u_{d}~dx>\int_{\Omega}K~dx.$ (1.3)
On the other hand, in [18], the authors justified that when $r(x)$ is constant, and therefore is independent of $K(x)$, the logistic equation with spatially heterogeneous resources always supports a total population strictly smaller than the total carrying capacity at equilibrium, which is just opposite to the case $r=\alpha K$.
DeAngelis et al [14] argued that, once the growth rate $r$ is not proportional to $K$, we can get different relations between the total population and the total carrying capacity. In particular, once these two functions are positively correlated (meaning that $r(x)=h(K(x))$ with $h^{\prime}>0$), the total population of $u_{d}$ exceeds the total carrying capacity, while for negatively correlated $r$ and $K$ ($r(x)=h(K(x))$ and $h^{\prime}<0$), the total population of $u_{d}$ is less for slow dispersal. The proofs of these results in [14] are strongly based on the estimation of the integral of $u_{d}$ for small $d>0$.
Unlike this local approach, in [25] inequality (1.3) was justified globally (and the opposite in [18] for the constant growth rate), for any $d>0$. In addition, it was noticed that the total solution equals the total of the carrying capacity when either $d=0$ or $d\to\infty$ when $r=\alpha K$. Then, a maximum of the average should be attained for some $d\in(0,\infty)$. Moreover, it was conjectured that
$M(d)=\int_{\Omega}u_{d}~dx$ (1.4)
is a unimodal function of $d$, increasing from $\int_{\Omega}K~dx$ to its maximum, and then decreasing, approaching the total carrying capacity from above for $d\to\infty$.
Connecting the results of [25] and [18], we notice that for $r=\alpha K=\alpha K^{1}$, the total population $u_{d}$ exceeds the total carrying capacity, while for $r=\alpha=\alpha K^{0}$ it is less. None of the two papers explores the case $r=\alpha K^{\lambda}$, $\lambda\in(0,1)$, and the current paper fills this gap. The naive suggestion that there is some critical $\lambda^{\ast}\in(0,1)$ such that for larger values (1.3) holds, while the opposite inequality is satisfied for smaller $\lambda$ is incorrect, as the situation is more complicated. Moreover, we consider any $\lambda$, not only in $[0,1]$.
In addition, we explore the model incorporating the third space-dependent parameter associated with dispersal. A strategy used to alleviate the effect of random diffusion with directed advection suggested in [3] assumed that we can model the diffusing quantity in (1.1) as not $u$, but $u/P$, where $P$ is the dispersal strategy adopted by the species. Then (1.2) is replaced by
$\begin{cases}\displaystyle d\Delta\left(\frac{u}{P}\right)+ru\left(1-\frac{u}{K}\right)=0,&x\in\Omega,\\ \displaystyle\frac{\partial}{\partial n}\left(\frac{u}{P}\right)=0,&x\in\partial\Omega.\end{cases}$ (1.5)
If $P$ is proportional to $K$, the carrying capacity is a solution of (1.5), independently of $d$, and species will evolve based on the available resources. Several weighted inequalities connecting $r,K,P$ with the stationary solution $u_{d}$ were obtained in [4, 22, 23] as auxiliary statements on the way of demonstrating that $P=K$ in (1.5) is a winning strategy, allowing the species choosing it either to protect its own habitat or invade a habitat of a competitor choosing a different way to disperse, or considering two different species cooperating in dispersal strategies in a sense that $K$ is a linear combination of two dispersal strategies with positive coefficients. However, explicit relations as in (1.3) have never been considered, and in this paper, we adapt the strategies from [14] to obtain the desired estimates. If several species coexist choosing different dispersal strategies, this can lead to coexistence when each of the populations occupies its niche in the spatial distribution of the available resources [1, 4]. If we relate the results of [16] to those in [25], this does not give an advantage to a competitor achieving higher density in the absence of others. Otherwise, an intermediate diffusion coefficient where the maximum density is attained would be optimal.
The purpose of this paper is two-fold:
  1. 1.
    We analyze the behaviour of the total population $M(d)$ when the intrinsic growth rate of species is described by $r=\alpha\left(\frac{K}{P}\right)^{\lambda}$. In particular, when $P\equiv 1$ as in (1.1), these results fill the gap between the results of [25] and [18].
  2. 2.
    For the model with three parameters $P,r,K$ as in (1.5), we explore the relation of the total population and the total carrying capacity, coinciding with (1.3) or the opposite inequality, and establish sufficient conditions for both. Overall, we investigate the dependency of the total population $M(d)$ on $d$;
In particular, we answer the following questions.
  1. 1.
    Are there specific choices of the dispersal strategy $P$ in (1.5) that guarantee (1.3)?
  2. 2.
    Since (1.3) holds for every $d\in(0,+\infty)$ when $r=\alpha\frac{K}{P}$, what happens when $r=\alpha\left(\frac{K}{P}\right)^{\lambda}$?
  3. 3.
    We demonstrate that, when $r=\alpha\frac{K}{P}$, the limits of $M(d)$ as $d\to 0^{+}$ and $d\to\infty$ are both equal to the integral of the carrying capacity. However, unlike the case of $P\equiv 1$ associated with the random (regular) diffusion, there are examples when $M(d)$ is not unimodal. Moreover, when the limits of $M(d)$ as $d\to 0^{+}$ and $d\to\infty$ are not equal, various forms of $M(d)$ are possible, from monotone to unimodal, or of a more complicated form.
Further, let us describe the structure of the paper. In Section 2 we justify the solution limits for infinitesimal or very large diffusion coefficients. Section 3 contains the main results of the paper. In particular, we justify under which conditions (1.5) is satisfied for any $d$, and explore the situations when either (1.5) or the opposite inequality are satisfied for small $d$. In Section 4, we provide examples illustrating the main results and simulating the cases when the theoretical results fail to establish the relations between the total population and the total carrying capacity. Section 5 contains discussion of both obtained results and remaining open questions. Appendix includes some additional weighted inequalities connecting the total population and the total carrying capacity of the environment.
2 Limit Population Sizes for Very Slow or Very Fast Diffusion
Let $\Omega$ be a bounded open region of $\mathbb{R}^{n}$. In this paper, we assume that
(A)   $r$, $K$, and $P$ are non-constant, positive parameters over $\overline{\Omega}$ such that $r$, $K$, $P\in C^{2}(\overline{\Omega})$.
Let us denote the Sobolev space $W^{1,2}(\Omega)=\left\{f\in L^{1}(\Omega):\,\,\,\nabla f\in L^{2}(\Omega)\right\}$, where $\nabla f$ is considered in the weak sense, equipped with the norm
$\left\|f\right\|_{W^{1,2}(\Omega)}=\left(\left\|f\right\|_{L^{2}(\Omega)}+\left\|\nabla f\right\|_{L^{2}(\Omega)}\right)^{\frac{1}{2}}.$
Denote by $u_{d}$ a solution for (1.5) for each $d>0$.
Lemma 2.1.
As $d\to 0^{+}$, we obtain
$u_{d}\to K\hskip 14.22636ptin\hskip 14.22636ptL^{\infty}(\overline{\Omega})\cap W^{1,2}(\Omega).$
Proof.
First we show convergence in $L^{\infty}(\Omega)$. For that purpose, we show that, for $d>0$ small, we obtain $\left\|u_{d}-K\right\|_{L^{\infty}(\Omega)}\leq\left(\alpha\max_{x\in\overline{\Omega}}P(x)\right)d$, where
$\alpha=2\frac{\left\|\Delta(\frac{K}{P})\right\|_{L^{\infty}(\overline{\Omega})}}{\min_{x\in\overline{\Omega}}(r(x)P(x))}.$
Denote the operator
$Lu=d\Delta\left(\frac{u}{P}\right)+ru\left(1-\frac{u}{K}\right).$
For each $d>0$, let $\overline{u}(x)=K(x)+\alpha P(x)d$, and compute
$L\overline{u}=$ $d\Delta\left(\frac{K}{P}\right)+d\Delta(\alpha d)+r(K+\alpha Pd)\left(-\frac{\alpha Pd}{K}\right)$
$=$ $d\left[\Delta\left(\frac{K}{P}\right)-\alpha rP\left(1+\frac{\alpha Pd}{K}\right)\right]$
$\leq$ $d\left[\Delta\left(\frac{K}{P}\right)-\alpha rP\right]$
$=$ $d\left[\Delta\left(\frac{K}{P}\right)-2\frac{\left\|\Delta(\frac{K}{P})\right\|_{L^{\infty}(\overline{\Omega})}}{\min_{x\in\overline{\Omega}}(r(x)P(x))}rP\right]$
$\leq$ $d\left[\Delta\left(\frac{K}{P}\right)-2\left\|\Delta\left(\frac{K}{P}\right)\right\|_{L^{\infty}(\overline{\Omega})}\right]\leq 0.$
Therefore, $\overline{u}$ is an upper solution for $Lu=0$.
If $\underline{u}(x)=K(x)-\alpha P(x)d$, we see that $\underline{u}>0$ once $d>0$ is small ($0<d<\min_{\overline{\Omega}}(K/P)/(2\alpha)$), and
$L\underline{u}=d\Delta\left(\frac{K}{P}\right)+r(K-\alpha Pd)\frac{\alpha Pd}{K}=d\left[\Delta\left(\frac{K}{P}\right)+\frac{1}{2}\alpha rP+\alpha rP\left(\frac{1}{2}-\frac{\alpha Pd}{K}\right)\right],$
and since $\frac{1}{2}-\frac{\alpha Pd}{K}>0$ for $0<d<\min_{\overline{\Omega}}(K/P)/(2\alpha)$, we obtain
$L\underline{u}$ $\geq d\left[\Delta\left(\frac{K}{P}\right)+\frac{1}{2}2\frac{\left\|\Delta(\frac{K}{P})\right\|_{L^{\infty}(\overline{\Omega})}}{\min_{x\in\overline{\Omega}}(r(x)P(x))}rP\right]$
$\geq d\left[\Delta\left(\frac{K}{P}\right)+\left\|\Delta\left(\frac{K}{P}\right)\right\|_{L^{\infty}(\overline{\Omega})}\right]\geq 0,$
and we conclude that $\underline{u}$ is a lower solution for $Lu=0$. And since $\underline{u}<\overline{u}$ in $\Omega$ and $u_{d}$ is the solution to $Lu=0$, we obtain $\underline{u}<u_{d}<\overline{u}$ in $\Omega$, which implies
$K(x)-\alpha P(x)d<u_{d}(x)<K(x)+\alpha P(x)d\implies|u_{d}(x)-K(x)|<\alpha P(x)d\leq\left(\alpha\max_{x\in\overline{\Omega}}P(x)\right)d,$
for any $x\in\Omega$ and $d>0$ small. Therefore, $u_{d}\to K$ in $L^{\infty}(\Omega)$ as $d\to 0^{+}$, and since $u_{d}$ is bounded and continuous in $\overline{\Omega}$ we also obtain convergence in $L^{\infty}(\overline{\Omega})$.
Finally, to show convergence on $W^{1,2}(\Omega)$, set $\varphi(x)=u_{d}-K(x)$ and notice that $\varphi$ satisfies
$d\Delta\left(\frac{\varphi}{P}\right)+d\Delta\left(\frac{K}{P}\right)+\frac{r}{K}u_{d}(-\varphi)=0.$ (2.1)
Multiplying (2.1) by $\frac{\varphi}{Pd}$ and integrating, we obtain
$\int_{\Omega}\Delta\left(\frac{\varphi}{P}\right)\frac{\varphi}{P}\,dx+\int_{\Omega}\Delta\left(\frac{K}{P}\right)\frac{\varphi}{P}\,dx-\int_{\Omega}\frac{r}{K}\frac{u_{d}\varphi^{2}}{Pd}\,dx=0$
$\implies$ $-\int_{\Omega}\left|\nabla\left(\frac{\varphi}{P}\right)\right|^{2}\,dx+\int_{\Omega}\Delta\left(\frac{K}{P}\right)\frac{\varphi}{P}\,dx-\int_{\Omega}\frac{r}{K}\frac{u_{d}\varphi^{2}}{Pd}\,dx=0$
and thus, we conclude that
$\int_{\Omega}\left|\nabla\left(\frac{\varphi}{P}\right)\right|^{2}\,dx=\int_{\Omega}\Delta\left(\frac{K}{P}\right)\frac{\varphi}{P}\,dx-\int_{\Omega}\frac{r}{K}\frac{u_{d}\varphi^{2}}{dP}\,dx\leq\int_{\Omega}\varphi\left\|\frac{1}{P}\Delta\left(\frac{K}{P}\right)\right\|_{L^{\infty}(\Omega)}\,dx,$
which tends to 0 as $d\to 0^{+}$, since $\varphi=u_{d}-K\to 0$ in $L^{\infty}(\Omega)$. Therefore, as $d\to 0^{+}$,
$\int_{\Omega}\left|\nabla\left(\frac{\varphi}{P}\right)\right|^{2}\,dx\to 0,$
and since $\frac{1}{P}$, $\nabla\left(\frac{1}{P}\right)$ are bounded on $\bar{\Omega}$, we conclude that $u_{d}\to K$ in $W^{1,2}(\Omega)$ as $d\to 0^{+}$. ∎
Lemma 2.2.
As $d\to+\infty$, we get
$u_{d}\to\beta P(x)\hskip 14.22636ptin\hskip 14.22636ptL^{\infty}(\overline{\Omega}),$
where $\beta=\frac{\int_{\Omega}rP\,dx}{\int_{\Omega}\frac{r}{K}P^{2}\,dx}$.
Proof.
Integrating (1.5) over $\Omega$ and using the boundary condition $\frac{\partial}{\partial n}(\frac{u_{d}}{P})=0$ on $\partial\Omega$, we obtain
$\int_{\Omega}ru_{d}\left(1-\frac{u_{d}}{K}\right)\,dx=0.$ (2.2)
Now, since $u_{d}$ and $\nabla u_{d}$ are bounded in $L^{\infty}(\Omega)$, we obtain from Arzelà-Ascoli theorem that there exists a subsequence $u_{d_{k}}$, $\{d_{k}\}\subset(0,+\infty)$ an increasing sequence, such that $u_{d_{k}}\to u_{\infty}$ uniformly in $\Omega$. Now, since from (1.5)
$\Delta\left(\frac{u_{d_{k}}}{P}\right)+\frac{1}{d_{k}}ru_{d_{k}}\left(1-\frac{u_{d_{k}}}{K}\right)=0$
for every $k\in\mathbb{N}$, take $k\to+\infty$ to obtain
$\Delta\left(\frac{u_{\infty}}{P}\right)=0,$
and since $\frac{\partial}{\partial n}(\frac{u_{\infty}}{P})=0$ in $\partial\Omega$, we obtain $u_{\infty}(x)=\beta P(x)$ for some constant $\beta\geq 0$.
From (2.2) applied to $d=d_{k}$, let $k\to+\infty$ to obtain
$0=\int_{\Omega}r(x)\beta P(x)\left(1-\frac{\beta P(x)}{K(x)}\right)\,dx=\beta\int_{\Omega}r(x)P(x)\,dx-\beta^{2}\int_{\Omega}\frac{r(x)}{K(x)}P(x)^{2}\,dx$
which implies
$\beta=0\,\,\,\,\textrm{ or }\,\,\,\,\beta=\frac{\int_{\Omega}rP\,dx}{\int_{\Omega}\frac{r}{K}P^{2}\,dx}.$
If we had $\beta=0$, then $u_{d_{k}}\to 0$ on the $L^{\infty}(\overline{\Omega})$ norm, for there exists $k\in\mathbb{N}$ such that $1-\frac{u_{d_{k}}}{K}>0$ in $\overline{\Omega}$. Since $u_{d_{k}}>0$ on $\overline{\Omega}$, we then have
$\int_{\Omega}ru_{d_{k}}\left(1-\frac{u_{d_{k}}}{K}\right)\,dx>0,$
which contradicts (2.2).
Corollary 2.3.
If $P=\alpha\frac{K}{r}$ for some $\alpha>0$ then
$\lim_{d\to+\infty}\int_{\Omega}u_{d}~dx=\int_{\Omega}Kdx.$
Proof.
From Proposition 2.2, $u_{d}\to\beta\frac{K}{r}$ in $L^{\infty}(\Omega)$ as $d\to+\infty$, where
$\beta=\frac{\int_{\Omega}r\frac{K}{r}\,dx}{\int_{\Omega}\frac{r}{K}\left[\frac{K}{r}\right]^{2}\,dx}=\frac{\int_{\Omega}K\,dx}{\int_{\Omega}\frac{K}{r}\,dx}.$
3 Main Results
3.1 The choice of the dispersal strategy to guarantee abundance
First of all, we explore the following question: can a species choose such a dispersal strategy $P$ that its total population exceeds the total carrying capacity? The answer is positive, once $P$ is proportional to the ratio of the carrying capacity and the intrinsic growth rate, while this intrinsic growth rate is non-constant. In other words, the areas of attraction should have higher available resources with smaller $r$.
Theorem 3.1.
Let (A) be satisfied and $P=\alpha\frac{K}{r}$ for some $\alpha>0$, then the solution $u_{d}$ of (1.5) satisfies (1.3) for every $d>0$.
Proof.
Setting $r=\alpha\frac{K}{P}$ on (1.5), we obtain
$\begin{cases}d\Delta\left(\frac{u_{d}}{P}\right)+\frac{\alpha}{P}u_{d}\left(K-u_{d}\right)=0&\mbox{on }\,\,\,\Omega,\\ \frac{\partial}{\partial n}\left(\frac{u_{d}}{P}\right)=0&\mbox{on }\,\partial\Omega.\end{cases}$ (3.1)
Multiply Equation (3.1) by $\frac{P}{u_{d}}$ and integrate over $\Omega$, to get
$d\int_{\Omega}\Delta\left(\frac{u_{d}}{P}\right)\frac{P}{u_{d}}\,dx+\alpha\int_{\Omega}(K-u_{d})\,dx=0.$ (3.2)
From the divergence theorem and the boundary condition of (3.1), we rewrite the first term of (3.2) as
$\int_{\Omega}\Delta\left(\frac{u_{d}}{P}\right)\frac{P}{u_{d}}\,dx$ $=-\int_{\Omega}\nabla\left(\frac{u_{d}}{P}\right)\cdot\nabla\left(\frac{P}{u_{d}}\right)\,dx=-\int_{\Omega}\nabla\left(\frac{u_{d}}{P}\right)\cdot\nabla\left(\left(\frac{u_{d}}{P}\right)^{-1}\right)\,dx$
$=\int_{\Omega}\nabla\left(\frac{u_{d}}{P}\right)\cdot\nabla\left(\frac{u_{d}}{P}\right)\left(\frac{u_{d}}{P}\right)^{-2}\,dx=\int_{\Omega}\left|\nabla\left(\frac{u_{d}}{P}\right)\right|^{2}\left(\frac{P}{u_{d}}\right)^{2}\,dx,$
and therefore, from (3.2),
$\alpha\int_{\Omega}(u_{d}-K)\,dx=d\int_{\Omega}\left|\nabla\left(\frac{u_{d}}{P}\right)\right|^{2}\left(\frac{P}{u_{d}}\right)^{2}\,dx>0.$
The last integral could only be zero if the ratio $\frac{u_{d}}{P}$ were constant over $\Omega$, leading to $u_{d}\equiv K$ being a positive solution to (3.1) which is also proportional to $P$. Then, the growth rate $r=\alpha\frac{K}{P}$ would be constant over $\Omega$, which we assume to be false. ∎
Remark 3.2.
In the case of $P=\alpha\frac{K}{r}$, it follows from Lemma 2.1 and from Corollary 2.3 that $\int_{\Omega}u_{d}\,dx$ converges to $\int_{\Omega}K\,dx$ when both $d\to 0^{+}$ and $d\to+\infty$. And since $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx$ for every $d>0$, the total population of solutions to (3.1) should be maximized for an intermediate diffusion coefficient $d^{*}\in(0,+\infty)$.
The assumption in (A) that $r$ is non-constant is quite significant: if $r$ is constant, the strategy chosen in Theorem 3.1 will lead to $P$ being proportional to $K$, leading to a solution identical to $K$ for any dispersal coefficient. And, as Theorem 3.3 below illustrates, this case maximizes all the possible population averages for a constant growth rate.
For the case where $P\equiv$ const, Lou [25] conjectures that $M(d)=\int_{\Omega}u_{d}\,dx$ should be a unimodal function. However, our analysis shows that for a general non-constant dispersal strategy $P(x)$, this unimodal behaviour might not hold, as illustrated in Example 4.3.
Note that the result in [25] is a particular case of Theorem 3.1 corresponding to $r$ being proportional to $K$ and $P\equiv 1$. Then, the winning strategy is free non-directed dispersal. If the intrinsic growth rate is not totally aligned with the carrying capacity, a choice of more sophisticated strategies is required.
The second question is what happens if the growth rate in the heterogeneous environment is constant.
Theorem 3.3.
Let $r>0$ be constant and $K$, $P>0$ be non-constant and linearly independent over $\Omega$. Then the solution $u_{d}$ of (1.5) satisfies
$\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx,~~\textrm{ for all }~~d>0.$
Proof.
The proof literally repeats the agrument in [18] for $P\equiv 1$, but it is presented here for completeness.
Assume the contrary that there exists some $d^{*}>0$ such that
$\int_{\Omega}u_{d^{*}}\,dx\geq\int_{\Omega}K\,dx.$ (3.3)
Integrating (1.5) for $d=d^{*}$, we get
$r\int_{\Omega}u_{d^{*}}\left(1-\frac{u_{d^{*}}}{K}\right)\,dx=-d^{*}\int_{\Omega}\Delta\left(\frac{u_{d^{*}}}{P}\right)\,dx=d^{*}\int_{\partial\Omega}\frac{\partial}{\partial n}\left(\frac{u_{d^{*}}}{P}\right)\,dS=0$
$\implies$ $\int_{\Omega}u_{d^{*}}\left(1-\frac{u_{d^{*}}}{K}\right)\,dx=0$ (3.4)
from the boundary condition in (1.5). Adding up (3.3) and (3.4), we obtain
$0\leq\int_{\Omega}\left(u_{d^{*}}-K+u_{d^{*}}\left(1-\frac{u_{d^{*}}}{K}\right)\right)\,dx=\int_{\Omega}\left(u_{d^{*}}-K\right)\left(1-\frac{u_{d^{*}}}{K}\right)\,dx=-\int_{\Omega}\frac{1}{K}\left(u_{d^{*}}-K\right)^{2}\,dx\leq 0,$
which leads to $\int_{\Omega}\frac{1}{K}\left(u_{d^{*}}-K\right)^{2}\,dx=0$ and $u_{d^{*}}\equiv K$ over $\Omega$, which is only possible if $P\equiv K$ over $\Omega$, contradicting the assumption that $P$ and $K$ are linearly independent.
Therefore, we must have $M(d)<\int_{\Omega}K\,dx$ for all $d>0$.
3.2 Tendencies for slow dispersal
In this section, we consider solution asymptotics in the case $d\to 0^{+}$.
Lemma 3.4.
Suppose that (A) holds and let $u_{d}$ be a solution to (1.5).
  1. 1.
    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$, then for $d>0$ small,
    $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx.$
  2. 2.
    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx<0$ , then for $d>0$ small,
    $\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx.$
Proof.
Let us prove item 1, as item 2 is proven analogously. Multiply the first equation in (1.5) by $\frac{K}{ru_{d}}$,
$d\Delta\left(\frac{u_{d}}{P}\right)\frac{K}{ru_{d}}+K-u_{d}=0,$
integrate it and use that $\frac{\partial}{\partial n}(\frac{u_{d}}{P})=0$ on $\partial\Omega$ to obtain
$d\int_{\Omega}\nabla\left(\frac{u_{d}}{P}\right)\cdot\nabla\left(\frac{K}{ru_{d}}\right)\,dx+\int_{\Omega}(K-u_{d})\,dx=0.$ (3.5)
Next, let us prove that
$\int_{\Omega}\left|\nabla\left(\frac{u_{d}}{P}\right)\cdot\nabla\left(\frac{K}{ru_{d}}\right)-\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\right|\,dx\to 0\textrm{ as }d\to 0^{+}.$
For that purpose, notice that
$\int_{\Omega}\left|\nabla\left(\frac{u_{d}}{P}\right)\cdot\nabla\left(\frac{K}{ru_{d}}\right)-\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\right|\,dx$
$\leq$ $\int_{\Omega}\left|\nabla\left(\frac{u_{d}}{P}-\frac{K}{P}\right)\cdot\nabla\left(\frac{K}{ru_{d}}\right)\right|\,dx+\int_{\Omega}\left|\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{K}{ru_{d}}-\frac{1}{r}\right)\right|\,dx$
$\leq$ $\left\|\nabla\left(\frac{u_{d}}{P}-\frac{K}{P}\right)\right\|_{L^{2}(\Omega)}\left\|\nabla\left(\frac{K}{ru_{d}}\right)\right\|_{L^{2}(\Omega)}+\left\|\nabla\left(\frac{K}{P}\right)\right\|_{L^{\infty}(\Omega)}\left\|\nabla\left(\frac{K}{ru_{d}}-\frac{1}{r}\right)\right\|_{L^{1}(\Omega)}$
$=$ $I+II.$
From Proposition 2.1 we get that $u_{d}\to K$ on $L^{\infty}(\Omega)\cap W^{1,2}(\Omega)$ as $d\to 0$, therefore the first term of $I$ vanishes. Moreover,
$\nabla\left(\frac{K}{ru_{d}}\right)=\nabla\left(\frac{K}{r}\right)\frac{1}{u_{d}}-\frac{K}{ru_{d}^{2}}\nabla u_{d},$
which is bounded, as $u_{d}\to K$ implies that $\nabla u_{d}$ is bounded and $u_{d}$ is far from zero almost everywhere. Therefore, $I\to 0$ when $d\to 0$.
To estimate $II$, notice that
$\left\|\nabla\left(\frac{K}{ru_{d}}-\frac{1}{r}\right)\right\|_{L^{1}(\Omega)}$
$=$ $\int_{\Omega}\left|\frac{\nabla K}{ru_{d}}-\frac{K\nabla r}{r^{2}u_{d}}-\frac{K}{ru_{d}^{2}}\nabla u_{d}+\frac{\nabla r}{r^{2}}\right|\,dx$
$\leq$ $\int_{\Omega}\left|\frac{\nabla K}{ru_{d}}-\frac{\nabla u_{d}}{ru_{d}}\right|\,dx+\int_{\Omega}\left|\frac{\nabla u_{d}}{ru_{d}}-\frac{K}{ru_{d}^{2}}\nabla u_{d}\right|\,dx+\int_{\Omega}\left|\frac{K\nabla r}{r^{2}u_{d}}-\frac{\nabla r}{r^{2}}\right|\,dx$
$=$ $\int_{\Omega}\left|\frac{1}{ru_{d}}\right|\left|\nabla K-\nabla u_{d}\right|\,dx+\int_{\Omega}\left|\frac{\nabla u_{d}}{ru_{d}}\right|\left|1-\frac{K}{u_{d}}\right|\,dx+\int_{\Omega}\left|\frac{\nabla r}{r^{2}}\right|\left|\frac{K}{u_{d}}-1\right|\,dx.$ (3.6)
According to Proposition 2.1, as $d\to 0^{+}$, $u_{d}\to K$ on both $L^{\infty}(\Omega)$ and $W^{1,2}(\Omega)$ norms, therefore, $|\frac{1}{u_{d}}|$ and $|\nabla u_{d}|$ are bounded on $\Omega$, and therefore, each integral on (3.6) vanishes as $d\to 0^{+}$.
Therefore, as $d\to 0^{+}$,
$\int_{\Omega}\nabla\left(\frac{u_{d}}{P}\right)\cdot\nabla\left(\frac{K}{ru_{d}}\right)\,dx=\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx+h(d),$
where $\lim_{d\to 0^{+}}h(d)=0$. And from (3.5), we get
$\int_{\Omega}(K-u_{d})\,dx=d\left(\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx+h(d)\right)$
which is negative for small enough $d>0$, since
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx=-\int_{\Omega}\frac{1}{r^{2}}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx\leq-\frac{1}{\max_{\bar{\Omega}}r^{2}}\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx<0.$
Remark 3.5.
When $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=0$, nothing can be concluded about the monotonicity of $\int_{\Omega}u_{d}\,dx$ for $d>0$ small, as shown in Example 4.1.
Definition 3.6.
We say that two functions $f:\Omega\to\mathbb{R}$ and $g:\Omega\to\mathbb{R}$ are correlated by a function $h:\mathbb{R}\to\mathbb{R}$ in $\Omega$ if we have
$f(x)=h(g(x))\mbox{ ~~ for all ~~ }x\in\Omega.$ (3.7)
If (3.7) is satisfied in $\Omega$ and $h$ is non-decreasing, we say that $f$ and $g$ are positively correlated in $\Omega$, and if $h$ is non-increasing, we say that $f$ and $g$ are negatively correlated in $\Omega$.
Corollary 3.7.
Suppose $r\equiv K$ over $\Omega$, and $P$ and $K$ are correlated by a function $h:(0,+\infty)\to(0,+\infty)$, and we write $P(x)=h(K(x))$ in $\Omega$.
  1. 1.
    If $h^{\prime}(t)>\frac{h(t)}{t}$ for all $t>0$, then for $d>0$ small,
    $\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx.$
  2. 2.
    If $h^{\prime}(t)<\frac{h(t)}{t}$ for all $t>0$, then for $d>0$ small,
    $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx.$
Proof.
  1. 1.
    By Lemma 3.4, it is enough to determine the sign of
    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx=\int_{\Omega}\left(\frac{|\nabla K|^{2}}{P}-K\frac{\nabla P\cdot\nabla K}{P^{2}}\right)\,dx.$
    If $P(x)=h(K(x))$, then
    $\nabla P=h^{\prime}(K(x))\nabla K(x),$
    and we obtain
    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx=\int_{\Omega}\left(P-Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}}\,dx=\int_{\Omega}\left(h(K)-Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}}\,dx.$
    If $h^{\prime}(t)>\frac{h(t)}{t}$ for all $t>0$, then $h(t)-th^{\prime}(t)<0$ for all $t>0$, and $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx<0$, which implies, by Lemma 3.4, that
    $\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx.$
    And if $h^{\prime}(t)<\frac{h(t)}{t}$ for all $t>0$, then $h(t)-th^{\prime}(t)>0$, implying $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx>0$, and follows from Lemma 3.4 that
    $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx.$
A sufficient condition for Lemma 3.4 is when there exists a positive or negative correlation between $r$ and $\frac{K}{P}$, a fact we prove next.
Theorem 3.8.
Assume assumption (A) holds. If $r$ and $\frac{K}{P}$ are positively correlated in $\Omega$, then, for $d>0$ small,
$\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx.$
If $r$ and $\frac{K}{P}$ are negatively correlated in $\Omega$, we get, for $d>0$ small,
$\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx.$
Proof.
If $r$ and $\frac{K}{P}$ are positively correlated (similar proof for negative correlation), there exists $h:\mathbb{R}\to\mathbb{R}$ non-decreasing with $r(x)=h\left(\frac{K(x)}{P(x)}\right)$ for every $x\in\Omega$, thus
$\nabla r=h^{\prime}\left(\frac{K}{P}\right)\nabla\left(\frac{K}{P}\right)$
and
$\nabla\left(\frac{K}{P}\right)\cdot\nabla r=h^{\prime}\left(\frac{K}{P}\right)\left|\nabla\left(\frac{K}{P}\right)\right|^{2}\geq 0\,\textrm{ everywhere in }\Omega.$
And since $h^{\prime}(x)\equiv 0$ or $\frac{K}{P}\equiv$ cte in $\Omega$ means that $r$ is constant, we obtain the strict inequality
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0.$
The proof is concluded by applying Lemma 3.4. ∎
Remark 3.9.
When $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$, we can only be sure that $M(d)=\displaystyle\int_{\Omega}u_{d}\,dx$ is maximized at some intermediate dispersion coefficient $d^{*}\in(0,+\infty)$ when $M(0)\geq M(+\infty)$, or equivalently,
$\int_{\Omega}K\,dx\geq\frac{\int_{\Omega}rP\,dx}{\int_{\Omega}\frac{r}{K}P^{2}\,dx}\int_{\Omega}P\,dx.$
If $M(0)<M(+\infty)$, it is possible for $M(d)$ to be always increasing over $(0,+\infty)$, as illustrated on Example 4.2.
Some additional relations of weighted averages are presented in the Appendix.
3.3 The case when the correlation is described by a power function
Based on Theorem 3.1, it is reasonable to question if solutions to (1.5) satisfy the estimate (1.3) when, instead of $r=\alpha\frac{K}{P}$, we have
$r=\alpha\left(\frac{K}{P}\right)^{\lambda}~~~\textrm{in}~~~\Omega,$ (3.8)
where $\alpha\in(0,+\infty)$ and $\lambda\in\mathbb{R}$. If $\lambda>0$, $r$ and $\frac{K}{P}$ are correlated by the function $h(x)=\alpha x^{\lambda}$, which is increasing in $(0,+\infty)$, and we can apply Theorem 3.8 to obtain an estimate (1.3) for small values of $d>0$.
In this section, let us denote by $u_{d,\lambda}$ the positive solution to
$\begin{cases}d\Delta\left(\frac{u(x)}{P(x)}\right)+\alpha\left(\frac{K(x)}{P(x)}\right)^{\lambda}u(x)\left(1-\frac{u(x)}{K(x)}\right)=0,&x\in\Omega,\\ \frac{\partial u}{\partial n}(x)=0,&x\in\partial\Omega,\end{cases}$ (3.9)
for each $\lambda\in\mathbb{R}$ and $d>0$, and the total population function
$M_{\lambda}(d)=\int_{\Omega}u_{d,\lambda}\,dx.$
To analyze the total population of $u_{d,\lambda}$ as $d$ increases, let us follow the strategy on Remark 3.9 and analyze the relation between $M_{\lambda}(0)$ and $M_{\lambda}(+\infty)$.
Lemma 3.10.
Let $f$, $g:\Omega\to\mathbb{R}$ be positive functions. If $f$ is non-constant and $\int_{\Omega}g\,dx=1$, it holds
$\frac{\int_{\Omega}gf\ln f\,dx}{\int_{\Omega}gf\,dx}>\int_{\Omega}g\ln f\,dx.$ (3.10)
Proof.
Notice that
$\int_{\Omega}g(x)f(x)\ln f(x)\,dx-\int_{\Omega}g(x)\ln f(x)\,dx\int_{\Omega}g(y)f(y)\,dy$
$=$ $\int_{\Omega}\left(f(x)-\int_{\Omega}g(y)f(y)\,dy\right)\left(\ln f(x)-\ln\left(\int_{\Omega}g(y)f(y)\,dy\right)\right)g(x)\,dx,$
since
$\int_{\Omega}\left(f(x)-\int_{\Omega}g(y)f(y)\,dy\right)\ln\left(\int_{\Omega}g(y)f(y)\,dy\right)g(x)\,dx=0,$
where we recall that $\int_{\Omega}g(x)\,dx=1$. Therefore, denoting $L=\int_{\Omega}g(y)f(y)\,dy$, we obtain
$\int_{\Omega}gf\ln f\,dx-\int_{\Omega}g\ln f\,dx\int_{\Omega}gf\,dy=\int_{\Omega}\left(f(x)-L\right)\left(\ln f(x)-\ln L\right)g(x)\,dx.$
We notice that this integral is non-negative, since $g(x)>0$, and $\ln y$ being increasing on $(0,+\infty)$ implies that f(x)Lf(x)-L and $\ln f(x)-\ln L$ always have the same signal over $\Omega$, and therefore,
$\left(f(x)-L\right)\left(\ln f(x)-\ln L\right)\geq 0~~~\textrm{for every}~~~x\in\Omega.$
Finally, $\int_{\Omega}\left(f(x)-L\right)\left(\ln f(x)-\ln L\right)g(x)\,dx=0$ if, and only if, $f(x)=L$ for all $x\in\Omega$, which contradicts $f$ being non-constant. ∎
Theorem 3.11.
The mapping $\lambda\mapsto M_{\lambda}(+\infty)=\frac{\int_{\Omega}\frac{K^{\lambda}}{P^{\lambda-1}}\,dx}{\int_{\Omega}\frac{K^{\lambda-1}}{P^{\lambda-2}}\,dx}\int_{\Omega}P\,dx$ is a strictly increasing function of $\lambda$ in $\mathbb{R}$.
Proof.
It is enough to show that
$F(\lambda)=\frac{\int_{\Omega}\frac{K^{\lambda}}{P^{\lambda-1}}\,dx}{\int_{\Omega}\frac{K^{\lambda-1}}{P^{\lambda-2}}\,dx}=\frac{\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-1}\,dx}{\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-2}\,dx}$ (3.11)
is strictly increasing in $\mathbb{R}$. Notice that
$\ln(F(\lambda))=\ln\left(\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-1}\,dx\right)-\ln\left(\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-2}\,dx\right),$
and differentiating with respect to $\lambda$,
$\frac{F^{\prime}(\lambda)}{F(\lambda)}=\frac{\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-1}\ln\left(\frac{K}{P}\right)\,dx}{\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-1}\,dx}-\frac{\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-2}\ln\left(\frac{K}{P}\right)\,dx}{\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-2}\,dx}.$ (3.12)
The right-hand side of Eq. (3.12) is positive, from (3.10) applied to
$f=\frac{K}{P},~~~~g=\left(\int_{\Omega}K\left(\frac{K}{P}\right)^{\lambda-2}\,dx\right)^{-1}K\left(\frac{K}{P}\right)^{\lambda-2}$
And since $F(\lambda)>0$, we obtain from (3.12) that $F^{\prime}(\lambda)>0$ in $\mathbb{R}$. Therefore, $F(\lambda)$ is increasing in $\mathbb{R}$. ∎
Theorem 3.11 and the fact that $M_{\lambda}(0)=M_{\lambda}(+\infty)$ for $\lambda=1$ implies that
$M_{\lambda}(0)>M_{\lambda}(+\infty)\textrm{ for }\lambda>1$
and
$M_{\lambda}(0)<M_{\lambda}(+\infty)\textrm{ for }\lambda<1.$
These facts are elucidated in Example 4.4, and we explore their implications for our understanding of the behaviour of $M_{\lambda}(d)$ in the Discussion section.
4 Examples
In this section, we illustrate some properties of the total population function $M(d)$ that were mentioned on Section 3. Note that, in this Section, all simulations were performed on the one-dimensional domain $\Omega=(0,1)$.
Example 4.1.
For small $d>0$, the monotonicity of the total population $M(d)=\int_{\Omega}u_{d}\,dx$ can be determined by Lemma 3.4 by evaluating the sign of $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx$. This first example illustrates that, when this integral vanishes, $M(d)$ can be either decreasing or increasing for small values of $d>0$.
Letting $K(x)=(\cos(2\pi x)+2)^{2}$, $P(x)=\sqrt{K(x)}$ and $r(x)=\cos(\pi x)+2$, we obtain
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx$ $=\int_{0}^{1}(\cos(2\pi x)+2)^{\prime}\cdot(\cos(\pi x)+2)^{\prime}\,dx=\int_{0}^{1}2\pi^{2}\sin(2\pi x)\sin(\pi x)\,dx$
$=\int_{0}^{1}4\pi^{2}\sin^{2}(\pi x)\cos(\pi x)\,dx=\frac{4\pi}{3}\sin^{3}(\pi x)\bigg|_{x=0}^{1}=0.$
And in this case, numerical implementations show that $\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx$ for $d>0$ is small, as seen in Figure 1.
Figure 1: Population decline when $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx=0$, for $K(x)=(\cos(2\pi x)+2)^{2}$, $P(x)=\sqrt{K(x)}$, and $r(x)=\cos(\pi x)+2$.
Now, taking $K(x)=(2x^{3}-3x^{2}+3)e^{2x^{3}-3x^{2}}$, $P(x)=e^{2x^{3}-3x^{2}}$ and $r(x)=\cos(2\pi x)+3$, we compute
$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=\int_{0}^{1}(2x^{3}-3x^{2}+3)^{\prime}\cdot(\cos(2\pi x)+3)^{\prime}\,dx=\int_{0}^{1}-12\pi(x^{2}-x)\sin(2\pi x)\,dx$
$=$ $6(x^{2}-x)\cos(2\pi x)\bigg|_{x=0}^{1}-\int_{0}^{1}6(2x-1)\cos(2\pi x)\,dx$
$=$ $0-\frac{3}{\pi}(2x-1)\sin(2\pi x)\bigg|_{x=0}^{1}+\int_{0}^{1}\frac{3}{\pi}2\sin(2\pi x)\,dx=0-\frac{3}{\pi^{2}}\cos(2\pi x)\bigg|_{x=0}^{1}=0.$
However, in this case, $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx$ for $d>0$ small, as seen in Figure 2.
Figure 2: Population growth when $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx=0$, for $K(x)=(2x^{3}-3x^{2}+3)e^{2x^{3}-3x^{2}}$, $P(x)=e^{2x^{3}-3x^{2}}$ and $r(x)=\cos(2\pi x)+3$.
Therefore, a more refined analysis is needed to determine the monotonicity of $M(d)$ when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=0$. For instance, if $r$ is constant on $\overline{\Omega}$, then [18] establishes that $M(d)<\int_{\Omega}K\,dx$ for all $d>0$, which implies that $M(d)$ should be decreasing near $d=0$.
Example 4.2.
This example explores the possible behaviours of $M(d)$ when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$. We know from Lemma 3.4 that $M(d)$ is increasing near $d=0$, and from Remark 3.9, a global maximum is attained at $d^{*}\in(0,+\infty)$ if $M(0)>M(+\infty)$. This scenario is illustrated in Figure 3a, which was generated using the parameters $K(x)=2+\cos(\pi x)$, $P(x)=1+\frac{1}{5}\cos(\pi x)$ and $r(x)=\frac{5}{4}+\frac{1}{4}\cos(\pi x)$.
It is also possible for $M(d)$ to be strictly increasing in $(0,+\infty)$, without attaining a local maximum, when $M(0)<M(+\infty)$. For instance, considering $K(x)=2+\cos(\pi x)$, $P(x)=1+\frac{1}{5}\cos(\pi x)$ and $r(x)=\exp(4\cos(\pi x))$, the total population increases monotonically with $d$, as shown in Figure 3b.
Figure 3: Different behaviours for the total population when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$, for $K(x)=2+\cos(\pi x)$, $P(x)=1+\frac{1}{5}\cos(\pi x)$ and (a) $r(x)=\frac{5}{4}+\frac{1}{4}\cos(\pi x)$, (b) $r(x)=\exp(4\cos(\pi x))$.
Example 4.3.
Returning to the case where $P$ and $\frac{K}{r}$ are proportional, recall from Remark 3.2 that $M(d)$ exceeds $\int_{\Omega}K\,dx$ for all $d>0$, with $M(d)$ approaching this integral as $d\to 0^{+}$ and $d\to+\infty$. In the case $P\equiv 1$, Lou [25] conjectured that $M(d)$ changes monotonicity only at the point $d^{*}\in(0,+\infty)$ where $M(d)$ attains its maximum.
We now demonstrate that non-constant dispersal strategies $P$ can invalidate this unimodality. In fact, let $K(x)=0.1+\cos(\pi x)+5\cos^{2}(\pi x)-2\cos^{3}(\pi x)$, $P(x)=1.5-3\cos(\pi x)+\cos^{2}(\pi x)+3\cos^{6}(\pi x)$ and $r(x)=\frac{K(x)}{P(x)}$. As shown in Figure 4, the corresponding total population $M(d)$ changes monotonicity three times and exhibits two different local maxima.
Figure 4: Non-unimodal total population curve under spatially heterogeneous dispersal strategy, for $K(x)=0.1+\cos(\pi x)+5\cos^{2}(\pi x)-2\cos^{3}(\pi x)$, $P(x)=1.5-3\cos(\pi x)+\cos^{2}(\pi x)+3\cos^{6}(\pi x)$ and $r(x)=\frac{K(x)}{P(x)}$.
Investigations are being conducted to examine the necessary conditions for the parameters $P$ and $K$ for which $M(d)$ changes monotonicity exactly once, which would be a key step toward proving Lou’s conjecture.
The last example illustrates the result of Theorem 3.11, which analyzes how the total population depends on the parameter $\lambda$ present in the growth rate $r(x)=\left(\frac{K(x)}{P(x)}\right)^{\lambda}$.
Example 4.4.
Let $K(x)=2+\cos(\pi x)$, $P(x)=2-\cos(2\pi x)$, and $r(x)=\left(\frac{K(x)}{P(x)}\right)^{\lambda}$. The parameter $\lambda$ is varied, taking the values $\lambda=-1$, 0, $0.5$, 11, $1.4$ and $2.3$ for comparison. Figure 5 shows the corresponding total population curves $M_{\lambda}(d)$.
Figure 5: Total population curves for $K(x)=2+\cos(\pi x)$, $P(x)=2-\cos(2\pi x)$, and $r(x)=\left(\frac{K(x)}{P(x)}\right)^{\lambda}$, where $\lambda=-1$, 0, $0.5$, 11, $1.4$ and $2.3$
First, with very slow diffusion, the same limit total population of $M_{\lambda}(0)=2$ is observed in every figure, while the fast dispersal limit $M_{\lambda}(+\infty)$ varies as an increasing function of $\lambda$, in agreement with Theorem 3.11. For $\lambda=-1$, we obtain negative correlation between $r$ and $\frac{K}{P}$, and $M(d)$ is a decreasing sigmoid-type function, which is in accordance with Theorem 3.8, and converges to $M_{\lambda}(+\infty)<M_{\lambda}(0)$ as $d\to+\infty$. The same behaviour of $M_{\lambda}(d)$ is observed for $\lambda=0$, leading to $M_{\lambda}(d)<\int_{\Omega}K\,dx$ for all $d>0$, as stated by [18].
For $\lambda=0.5$, the function $M_{\lambda}(d)$ takes a unimodal form, initially increasing due to the positive correlation between $r$ and $\frac{K}{P}$ (Theorem 3.8), reaching a maximum value, and eventually decreasing to a value below $M_{\lambda}(0)$. For $\lambda=1$, following Lou’s conjecture, $M_{\lambda}(d)$ is unimodal with equal limit values as $d\to 0^{+}$ and $d\to+\infty$, thereby validating Theorem 3.1 and Remark 3.2.
For $\lambda>1$, $r$ and $\frac{K}{P}$ are positively correlated, and Theorem 3.11 guarantees that the limit value $M_{\lambda}(+\infty)$ exceeds $M_{\lambda}(0^{+})$. The transient behaviour, however, differs for $\lambda=1.4$ and $\lambda=2.3$, as for the first case $M_{\lambda}(d)$ is unimodal, attaining a maximum value in $(0,+\infty)$, whereas for $\lambda=2.3$, $M_{\lambda}(d)$ takes the monotone increasing sigmoid form.
Lastly, comparing all simulations, we observe the maximum and minimum population levels are non-decreasing functions of $\lambda$. Further analysis is underway to determine if this relation holds universally.
5 Discussion
We start with briefly summarizing the findings of the paper.
  1. 1.
    We identified the case when diffusion leads to population increase, for any of its values. As a special case for $P\equiv 1$, this includes Lou’s result in [25].
  2. 2.
    Assuming positive (or negative) correlation between the growth rate and the ratio of the carrying capacity to the dispersal strategy, which for high levels of diffusion characterizes per capita space-dependent available resources, we determine when a small diffusion has a positive (negative) effect on the total population size compared to its immobile counterpart.
  3. 3.
    In the particular case when the growth rate is a power function of the ratio $r=(K/P)^{\lambda}$, we can characterize the global population size for large disparsal speed. We conclude that, once the species adopt a large enough diffusion rate, the total population exceeds the total of the carrying capacity, once the power function is convex, and is less if the function is concave. This fact has never been established before for the logistic equation with a regular diffusion. The two cases that have been already investigated in the particular case $P\equiv 1$ are $\lambda=1$ [25] and $\lambda=0$ [18]. We illustrate that the values $\lambda=0$ and $\lambda=1$ are in some sense critical.
Let us comment on the third item. Note that the intermediate behaviour of the total population size as a function of the diffusion coefficient exhibits distinct profiles depending on the value of $\lambda$, as illustrated in Fig. 5. Based on the results of this work, we conclude that for the case where $r=(K/P)^{\lambda}$, the following holds:
  1. 1.
    For $\lambda\leq 0$, the average population decreases from its carrying capacity at $d=0$. For $\lambda<0$, this is justified from $r=\alpha\left(\frac{K}{P}\right)^{\lambda}$ representing a negative correlation between $r$ and $\frac{K}{P}$ and from Theorem 3.8, while the $\lambda=0$ case is covered by [18]. Furthermore, since Theorem 3.11 gives $M_{\lambda}(+\infty)<M_{\lambda}(0)$, we expect that $M_{\lambda}(d)<M_{\lambda}(0)$ for all $d\in(0,+\infty)$.
  2. 2.
    For $0<\lambda<1$, $r=\alpha\left(\frac{K}{P}\right)^{\lambda}$ leads to a positive correlation between $r$ and $\frac{K}{P}$, and Theorem 3.8 implies that the average population exceeds the average carrying capacity for slow diffusion $d>0$. However, the inequality $M_{\lambda}(d)>\int_{\Omega}K\,dx$ cannot be satisfied for all $d>0$, specifically for $d$ large enough, since the total population of $u_{d,\lambda}$ converges to $M_{\lambda}(+\infty)<\int_{\Omega}K\,dx$ as $d\to+\infty$. Therefore, $M_{\lambda}(d)$ is maximized for some diffusion coefficient $d^{*}\in(0,+\infty)$.
  3. 3.
    For $\lambda=1$, Theorem 3.1 asserts that the total population exceeds the total carrying capacity for every $d>0$. Due to the continuity of solutions to (3.9) with respect to $\lambda$, we expect that for values of $\lambda$ close to $1$, the average population $M_{\lambda}(d)$ is still maximized at some intermediate point $d_{\lambda}^{*}\in(0,+\infty)$.
  4. 4.
    For $\lambda>1$ large enough, the asymptotic value $M_{\lambda}(+\infty)$ should also become large enough that $M_{\lambda}(d)$ is an increasing function of $d$ in $(0,+\infty)$.
Based on the preceding analysis, we formulate the following conjecture:
Conjecture.
Assume $K$, $P$ and $r=\alpha\left(\frac{K}{P}\right)^{\lambda}$ satisfy hypothesis (A). Then there exists a critical value $\lambda^{*}>1$ such that
  • a)
    If $\lambda\in(0,\lambda^{*})$, the function $M_{\lambda}(d)$ attains its maximum at a finite diffusion rate $d^{*}_{\lambda}\in(0,+\infty)$;
  • b)
    If $\lambda\in(\lambda^{*},+\infty)$, the function $M_{\lambda}(d)$ is strictly increasing in $(0,+\infty)$, and thus does not attain a maximum value.
With these empirical results and some statements justified, especially in the asymptotic cases of very slow or very fast dispersal, the problem of the qualitative description of the population size as a function of the diffusion coefficient still leaves many questions without answer. Let us discuss some topics emerging from the research of the current paper.
  1. 1.
    The first group of questions, as mentioned above, is related to the form of the total population as a function of the diffusion coefficient.
    1. (a)
      Investigation of the case $r=f(K/P)$, where $f(x)=x^{\lambda}$, opens ways to the study of the dependency of the population on the diffusion coefficient for $f(x)=ax+b$, fractional linear, exponential (in this case, we expect the behaviour similar to $f(x)=x^{\lambda}$ with large positive $\lambda$) etc.
    2. (b)
      Prove or disprove for $P\equiv 1$ the conjecture that for $r$ proportional to $K$, the total population is a unimodal function in $d$.
    3. (c)
      If the previous item is justified, what are the conditions on $P,r,K$ (other than constant $P$ and proportional $r,K$) allowing to conclude that the total population is a unimodal function in $d$?
    4. (d)
      What are the relations of $P,r,K$ leading to the monotone dependency of the total population on the diffusion coefficient?
    5. (e)
      Prove or disprove the conjecture that, once $r=f(K/P)$, where $f$ is an increasing function, the total population for large diffusion exceeds the total carrying capacity for the convex $f$ and is less for concave $f$. Note that the fact of equality for linear $h$ follows from Corollary 2.3.
  2. 2.
    Maximizing the total population size is not directly connected to the success in the compertition, or evolution stability, compare, for example, [5, 22, 23, 25]. However, there is a process which is closely connected to the population size - harvesting. Consider for the model
    $\begin{cases}d\Delta\left(\frac{u}{P}\right)+ru\left(1-\frac{u}{K}\right)-Eu=0,&x\in\Omega,\\ \frac{\partial}{\partial n}\left(\frac{u}{P}\right)=0,&x\in\partial\Omega\end{cases}$ (5.1)
    the problem of maximizing the Maximum Sustainable Yield (MSY) $\int_{\Omega}E(x)u(x)~dx$. In [3] for $P\equiv K$, it was noticed that $E=r/2$ leads to MSY. With some non-constant $P\not\equiv K$, for which $r$ is MSY achieved? If, in addition to $E$, we can control the diffusion coefficient $d$, how is MSY $d$-dependent?
  3. 3.
    We explored the dependency of the average solution on $\lambda$ when $r=(K/P)^{\lambda}$. Let us contemplate whether some of the observations in numerical examples can be justified theoretically.
    1. (a)
      Is the maximum population value increasing as the function of $\lambda$, not only the limit population value for very fast dispersals?
    2. (b)
      There is analysis of the limit $M_{\lambda}(+\infty)$. Has this function a limit for $\lambda\to+\infty$? The same question can be considered for the supremum value of the population. However, if the conjecture is justified, for large $\lambda$, the value of $M_{\lambda}(+\infty)$ is also the supremum value.
    3. (c)
      In our analysis, we fixed $P$ and $K$ and got variable $r$. Is there any differece if $r,K$ are fixed, and we modify $P=K/r^{1/\lambda}$?
Acknowledgment
The authors were partially supported by the NSERC Discovery Grant # RGPIN-2020-03934.

### v2

@@
 Differential Equ.), Lou proved that, once the intrinsic growth rate $r$ in the logistic model is proportional to the spatially heterogeneous carrying capacity $K$ ($r=K^{1}$), the total population under the regular diffusion exceeds the total of the carrying capacity.
-He also conjectured that the dependency of the total population on the diffusion coefficient is unimodal, increasing to its maximum and then decreasing to the asymptote which is the total of the carrying capacity.
 DeAngelis et al (J.
@@
 This connection was explored in [14, 15, 18, 26], where population dynamics was modelled using the logistic equation with random diffusion
-$\begin{cases}\frac{\partial u}{\partial t}(x,t)=d\Delta u(t,x)+r(x)u(t,x)\left(1-\frac{u(t,x)}{K(x)}\right)=0,&x\in\Omega,~~t>0\\ u(x,0)=u_{0}(x),&x\in\Omega\\ \frac{\partial u}{\partial n}(t,x)=0,&x\in\partial\Omega,~~t>0.\end{cases}$ (1.1)
+$\begin{cases}\frac{\partial u}{\partial t}(x,t)=d\Delta u(t,x)+r(x)u(t,x)\left(1-\frac{u(t,x)}{K(x)}\right),&x\in\Omega,~~t>0\\ u(x,0)=u_{0}(x),&x\in\Omega\\ \frac{\partial u}{\partial n}(t,x)=0,&x\in\partial\Omega,~~t>0.\end{cases}$ (1.1)
 Here $r(x)$ and $K(x)$ are time-independent positive on $\bar{\Omega}$ functions representing the intrinsic growth rate and the carrying capacity, respectively.
@@
 On the other hand, in [18], the authors justified that when $r(x)$ is constant, and therefore is independent of $K(x)$, the logistic equation with spatially heterogeneous resources always supports a total population strictly smaller than the total carrying capacity at equilibrium, which is just opposite to the case $r=\alpha K$.
+This result can be interpreted as, once the growth rate of species is unchanged over the domain, a uniform distribution of resources maximizes the population of species.
+To strengthen this result, Mazari-Fouquer [28] quantifies the optimality of a constant carrying capacity on the maximization of average population of species, compared to other resource distribution profiles.
 DeAngelis et al [14] argued that, once the growth rate $r$ is not proportional to $K$, we can get different relations between the total population and the total carrying capacity.
@@
 Then, a maximum of the average should be attained for some $d\in(0,\infty)$.
-Moreover, it was conjectured that
+Although [26] conjectures that
 $M(d)=\int_{\Omega}u_{d}~dx$ (1.4)
-is a unimodal function of $d$, increasing from $\int_{\Omega}K~dx$ to its maximum, and then decreasing, approaching the total carrying capacity from above for $d\to\infty$.
+attains exactly one local maximum in $(0,+\infty)$, it was proven by [25] the existence of a carrying capacity profile that induces multiple local maxima for the average of $u_{d}$.
 Connecting the results of [26] and [18], we notice that for $r=\alpha K=\alpha K^{1}$, the total population $u_{d}$ exceeds the total carrying capacity, while for $r=\alpha=\alpha K^{0}$ it is less.
@@
     Since (1.3) holds for every $d\in(0,+\infty)$ when $r=\alpha\frac{K}{P}$, what happens when $r=\alpha\left(\frac{K}{P}\right)^{\lambda}$?
-  3. 3.
-    We demonstrate that, when $r=\alpha\frac{K}{P}$, the limits of $M(d)$ as $d\to 0^{+}$ and $d\to\infty$ are both equal to the integral of the carrying capacity.
-However, unlike the case of $P\equiv 1$ associated with the random (regular) diffusion, there are examples when $M(d)$ is not unimodal.
-Moreover, when the limits of $M(d)$ as $d\to 0^{+}$ and $d\to\infty$ are not equal, various forms of $M(d)$ are possible, from monotone to unimodal, or of a more complicated form.
 Further, let us describe the structure of the paper.
@@
 Let $\Omega$ be a bounded open region of $\mathbb{R}^{n}$. In this paper, we assume that
-(A)   $r$, $K$, and $P$ are non-constant, positive parameters over $\overline{\Omega}$ such that $r$, $K$, $P\in C^{2}(\overline{\Omega})$.
+(A)   $r$, $K$, and $P$ are non-constant, positive functions over $\overline{\Omega}$ such that $r$, $K$, $P\in C^{2}(\overline{\Omega})$.
 Let us denote the Sobolev space $W^{1,2}(\Omega)=\left\{f\in L^{1}(\Omega):\,\,\,\nabla f\in L^{2}(\Omega)\right\}$, where $\nabla f$ is considered in the weak sense, equipped with the norm
@@
 And since $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx$ for every $d>0$, the total population of solutions to (3.1) should be maximized for an intermediate diffusion coefficient $d^{*}\in(0,+\infty)$.
+However, multiple local maxima for $\int_{\Omega}u_{d}\,dx$ might exist, as illustrated in Example 4.2.
 The assumption in (A) that $r$ is non-constant is quite significant: if $r$ is constant, the strategy chosen in Theorem 3.1 will lead to $P$ being proportional to $K$, leading to a solution identical to $K$ for any dispersal coefficient.
 And, as Theorem 3.3 below illustrates, this case maximizes all the possible population averages for a constant growth rate.
-For the case where $P\equiv$ const, Lou [25] conjectures that $M(d)=\int_{\Omega}u_{d}\,dx$ should be a unimodal function.
-However, our analysis shows that for a general non-constant dispersal strategy $P(x)$, this unimodal behaviour might not hold, as illustrated in Example 4.3.
 Note that the result in [26] is a particular case of Theorem 3.1 corresponding to $r$ being proportional to $K$ and $P\equiv 1$.
@@
 Proof.
-The proof literally repeats the agrument in [18] for $P\equiv 1$, but it is presented here for completeness.
+The proof literally repeats the argument in [18] for $P\equiv 1$, but it is presented here for completeness.
 Assume the contrary that there exists some $d^{*}>0$ such that
@@
   1. 1.
-    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$, then for $d>0$ small,
+    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0$, then for $d>0$ small,
     $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx.$
   2. 2.
-    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx<0$ , then for $d>0$ small,
+    If $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx>0$ , then for $d>0$ small,
     $\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx.$
@@
 $\int_{\Omega}(K-u_{d})\,dx=d\left(\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx+h(d)\right)$
-which is negative for small enough $d>0$, since
-$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx=-\int_{\Omega}\frac{1}{r^{2}}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx\leq-\frac{1}{\max_{\bar{\Omega}}r^{2}}\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx<0.$
-Remark 3.5.
-When $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=0$, nothing can be concluded about the monotonicity of $\int_{\Omega}u_{d}\,dx$ for $d>0$ small, as shown in Example 4.1.
+which is negative for small enough $d>0$, since $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0$.
 Definition 3.5.
@@
     By Lemma 3.4, it is enough to determine the sign of
-    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx=\int_{\Omega}\left(\frac{|\nabla K|^{2}}{P}-K\frac{\nabla P\cdot\nabla K}{P^{2}}\right)\,dx.$
+    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{K}\right)\,dx=\int_{\Omega}\left(-\frac{|\nabla K|^{2}}{PK^{2}}+K\frac{\nabla P\cdot\nabla K}{P^{2}K^{2}}\right)\,dx.$
     If $P(x)=h(K(x))$, then
@@
     and we obtain
-    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx=\int_{\Omega}\left(P-Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}}\,dx=\int_{\Omega}\left(h(K)-Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}}\,dx.$
-    If $h^{\prime}(t)>\frac{h(t)}{t}$ for all $t>0$, then $h(t)-th^{\prime}(t)<0$ for all $t>0$, and $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx<0$, which implies, by Lemma 3.4, that
+    $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{K}\right)\,dx=\int_{\Omega}\left(-P+Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}K^{2}}\,dx=\int_{\Omega}\left(-h(K)+Kh^{\prime}(K)\right)\frac{|\nabla K|^{2}}{P^{2}K^{2}}\,dx.$
+    If $h^{\prime}(t)>\frac{h(t)}{t}$ for all $t>0$, then $h(t)+th^{\prime}(t)>0$ for all $t>0$, and $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{K}\right)\,dx>0$, which implies, by Lemma 3.4, that
     $\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx.$
-    And if $h^{\prime}(t)<\frac{h(t)}{t}$ for all $t>0$, then $h(t)-th^{\prime}(t)>0$, implying $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla K\,dx>0$, and follows from Lemma 3.4 that
+    And if $h^{\prime}(t)<\frac{h(t)}{t}$ for all $t>0$, then $h(t)+th^{\prime}(t)<0$, implying $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{K}\right)\,dx<0$, and follows from Lemma 3.4 that
     $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx.$
@@
 If $r$ and $\frac{K}{P}$ are positively correlated (similar proof for negative correlation), there exists $h:\mathbb{R}\to\mathbb{R}$ non-decreasing with $r(x)=h\left(\frac{K(x)}{P(x)}\right)$ for every $x\in\Omega$, thus
-$\nabla r=h^{\prime}\left(\frac{K}{P}\right)\nabla\left(\frac{K}{P}\right)$
+$\nabla\left(\frac{1}{r}\right)=-h\left(\frac{K}{P}\right)^{-2}h^{\prime}\left(\frac{K}{P}\right)\nabla\left(\frac{K}{P}\right)$
 and
-$\nabla\left(\frac{K}{P}\right)\cdot\nabla r=h^{\prime}\left(\frac{K}{P}\right)\left|\nabla\left(\frac{K}{P}\right)\right|^{2}\geq 0\,\textrm{ everywhere in }\Omega.$
+$\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)=-\frac{h^{\prime}\left(K/P\right)}{h\left(K/P\right)^{2}}\left|\nabla\left(\frac{K}{P}\right)\right|^{2}\leq 0\,\textrm{ everywhere in }\Omega.$
 And since $h^{\prime}(x)\equiv 0$ or $\frac{K}{P}\equiv$ cte in $\Omega$ means that $r$ is constant, we obtain the strict inequality
-$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0.$
+$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0.$
 The proof is concluded by applying Lemma 3.4. ∎
 Remark 3.8.
-When $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$, we can only be sure that $M(d)=\displaystyle\int_{\Omega}u_{d}\,dx$ is maximized at some intermediate dispersion coefficient $d^{*}\in(0,+\infty)$ when $M(0)\geq M(+\infty)$, or equivalently,
+When $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0$, we can only be sure that $M(d)=\displaystyle\int_{\Omega}u_{d}\,dx$ is maximized at some intermediate dispersion coefficient $d^{*}\in(0,+\infty)$ when $M(0)\geq M(+\infty)$, or equivalently,
 $\int_{\Omega}K\,dx\geq\frac{\int_{\Omega}rP\,dx}{\int_{\Omega}\frac{r}{K}P^{2}\,dx}\int_{\Omega}P\,dx.$
@@
 $M_{\lambda}(d)=\int_{\Omega}u_{d,\lambda}\,dx.$
-To analyze the total population of $u_{d,\lambda}$ as $d$ increases, let us follow the strategy on Remark 3.9 and analyze the relation between $M_{\lambda}(0)$ and $M_{\lambda}(+\infty)$.
+To analyze the total population of $u_{d,\lambda}$ as $d$ increases, let us follow the strategy of Remark 3.8 and analyze the relation between $M_{\lambda}(0)$ and $M_{\lambda}(+\infty)$.
 Lemma 3.9.
@@
 Example 4.1.
-For small $d>0$, the monotonicity of the total population $M(d)=\int_{\Omega}u_{d}\,dx$ can be determined by Lemma 3.4 by evaluating the sign of $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx$.
-This first example illustrates that, when this integral vanishes, $M(d)$ can be either decreasing or increasing for small values of $d>0$.
-Letting $K(x)=(\cos(2\pi x)+2)^{2}$, $P(x)=\sqrt{K(x)}$ and $r(x)=\cos(\pi x)+2$, we obtain
-$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx$ $=\int_{0}^{1}(\cos(2\pi x)+2)^{\prime}\cdot(\cos(\pi x)+2)^{\prime}\,dx=\int_{0}^{1}2\pi^{2}\sin(2\pi x)\sin(\pi x)\,dx$
-$=\int_{0}^{1}4\pi^{2}\sin^{2}(\pi x)\cos(\pi x)\,dx=\frac{4\pi}{3}\sin^{3}(\pi x)\bigg|_{x=0}^{1}=0.$
-And in this case, numerical implementations show that $\int_{\Omega}u_{d}\,dx<\int_{\Omega}K\,dx$ for $d>0$ is small, as seen in Figure 1.
-Figure 1: Population decline when $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx=0$, for $K(x)=(\cos(2\pi x)+2)^{2}$, $P(x)=\sqrt{K(x)}$, and $r(x)=\cos(\pi x)+2$.
-Now, taking $K(x)=(2x^{3}-3x^{2}+3)e^{2x^{3}-3x^{2}}$, $P(x)=e^{2x^{3}-3x^{2}}$ and $r(x)=\cos(2\pi x)+3$, we compute
-$\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=\int_{0}^{1}(2x^{3}-3x^{2}+3)^{\prime}\cdot(\cos(2\pi x)+3)^{\prime}\,dx=\int_{0}^{1}-12\pi(x^{2}-x)\sin(2\pi x)\,dx$
-$=$ $6(x^{2}-x)\cos(2\pi x)\bigg|_{x=0}^{1}-\int_{0}^{1}6(2x-1)\cos(2\pi x)\,dx$
-$=$ $0-\frac{3}{\pi}(2x-1)\sin(2\pi x)\bigg|_{x=0}^{1}+\int_{0}^{1}\frac{3}{\pi}2\sin(2\pi x)\,dx=0-\frac{3}{\pi^{2}}\cos(2\pi x)\bigg|_{x=0}^{1}=0.$
-However, in this case, $\int_{\Omega}u_{d}\,dx>\int_{\Omega}K\,dx$ for $d>0$ small, as seen in Figure 2.
-Figure 2: Population growth when $\int_{\Omega}\nabla(K/P)\cdot\nabla r\,dx=0$, for $K(x)=(2x^{3}-3x^{2}+3)e^{2x^{3}-3x^{2}}$, $P(x)=e^{2x^{3}-3x^{2}}$ and $r(x)=\cos(2\pi x)+3$.
-Therefore, a more refined analysis is needed to determine the monotonicity of $M(d)$ when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx=0$.
-For instance, if $r$ is constant on $\overline{\Omega}$, then [18] establishes that $M(d)<\int_{\Omega}K\,dx$ for all $d>0$, which implies that $M(d)$ should be decreasing near $d=0$.
-Example 4.2.
-This example explores the possible behaviours of $M(d)$ when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla r\,dx>0$.
+This example explores the possible behaviours of $M(d)$ when $\int_{\Omega}\nabla\left(\frac{K}{P}\right)\cdot\nabla\left(\frac{1}{r}\right)\,dx<0$.
 We know from Lemma 3.4 that $M(d)$ is increasing near $d=0$, and from Remark 3.8, a global maximum is attained at $d^{*}\in(0,+\infty)$ if $M(0)>M(+\infty)$.
@@
 Example 4.2.
-Returning to the case where $P$ and $\frac{K}{r}$ are proportional, recall from Remark 3.2 that $M(d)$ exceeds $\int_{\Omega}K\,dx$ for all $d>0$, with $M(d)$ approaching this integral as $d\to 0^{+}$ and $d\to+\infty$.
-In the case $P\equiv 1$, Lou [25] conjectured that $M(d)$ changes monotonicity only at the point $d^{*}\in(0,+\infty)$ where $M(d)$ attains its maximum.
-We now demonstrate that non-constant dispersal strategies $P$ can invalidate this unimodality.
-In fact, let $K(x)=0.1+\cos(\pi x)+5\cos^{2}(\pi x)-2\cos^{3}(\pi x)$, $P(x)=1.5-3\cos(\pi x)+\cos^{2}(\pi x)+3\cos^{6}(\pi x)$ and $r(x)=\frac{K(x)}{P(x)}$.
-As shown in Figure 4, the corresponding total population $M(d)$ changes monotonicity three times and exhibits two different local maxima.
+Returning to the case where $P$ and $\frac{K}{r}$ are proportional, recall from Remark 3.2 that $M(d)$ exceeds $\int_{\Omega}K\,dx$ for all $d>0$, with $M(d)$ approaching this integral as $d\to 0^{+}$ and $d\to+\infty$, implying that $M(d)$ attains a maximum value at $d^{*}\in(0,+\infty)$.
+We now illustrate the local maximum value of $M(d)$ my not be unique.
+Let $K(x)=0.1+\cos(\pi x)+5\cos^{2}(\pi x)-2\cos^{3}(\pi x)$, $P(x)=1.5-3\cos(\pi x)+\cos^{2}(\pi x)+3\cos^{6}(\pi x)$ and $r(x)=\frac{K(x)}{P(x)}$.
+As shown in Figure 2, the corresponding total population $M(d)$ exhibits two different local maxima.
 Figure 2: Non-unimodal total population curve under spatially heterogeneous dispersal strategy, for $K(x)=0.1+\cos(\pi x)+5\cos^{2}(\pi x)-2\cos^{3}(\pi x)$, $P(x)=1.5-3\cos(\pi x)+\cos^{2}(\pi x)+3\cos^{6}(\pi x)$ and $r(x)=\frac{K(x)}{P(x)}$.
-Investigations are being conducted to examine the necessary conditions for the parameters $P$ and $K$ for which $M(d)$ changes monotonicity exactly once, which would be a key step toward proving Lou’s conjecture.
+Investigations are being conducted to examine the necessary conditions for the parameters $P$ and $K$ for which $M(d)$ changes monotonicity exactly once and attains one unique maximum value.
 The last example illustrates the result of Theorem 3.10, which analyzes how the total population depends on the parameter $\lambda$ present in the growth rate $r(x)=\left(\frac{K(x)}{P(x)}\right)^{\lambda}$.
@@
 For $\lambda=0.5$, the function $M_{\lambda}(d)$ takes a unimodal form, initially increasing due to the positive correlation between $r$ and $\frac{K}{P}$ (Theorem 3.7), reaching a maximum value, and eventually decreasing to a value below $M_{\lambda}(0)$.
-For $\lambda=1$, following Lou’s conjecture, $M_{\lambda}(d)$ is unimodal with equal limit values as $d\to 0^{+}$ and $d\to+\infty$, thereby validating Theorem 3.1 and Remark 3.2.
+For $\lambda=1$, $M_{\lambda}(d)$ is always greater than $\int_{\Omega}K\,dx$ with equal limit values as $d\to 0^{+}$ and $d\to+\infty$, thereby validating Theorem 3.1 and Remark 3.2.
 For $\lambda>1$, $r$ and $\frac{K}{P}$ are positively correlated, and Theorem 3.10 guarantees that the limit value $M_{\lambda}(+\infty)$ exceeds $M_{\lambda}(0^{+})$.
@@
   3. 3.
-    In the particular case when the growth rate is a power function of the ratio $r=(K/P)^{\lambda}$, we can characterize the global population size for large disparsal speed.
+    In the particular case when the growth rate is a power function of the ratio $r=(K/P)^{\lambda}$, we can characterize the global population size for large dispersal speed.
 We conclude that, once the species adopt a large enough diffusion rate, the total population exceeds the total of the carrying capacity, once the power function is convex, and is less if the function is concave.
@@
 Note that the intermediate behaviour of the total population size as a function of the diffusion coefficient exhibits distinct profiles depending on the value of $\lambda$, as illustrated in Fig.
-5.
+3.
 Based on the results of this work, we conclude that for the case where $r=(K/P)^{\lambda}$, the following holds:
@@
     2. (b)
-      Prove or disprove for $P\equiv 1$ the conjecture that for $r$ proportional to $K$, the total population is a unimodal function in $d$.
+      What are the relations of $P,r,K$ leading to the monotone or the unimodal dependency of the total population on the diffusion coefficient?
     3. (c)
-      If the previous item is justified, what are the conditions on $P,r,K$ (other than constant $P$ and proportional $r,K$) allowing to conclude that the total population is a unimodal function in $d$?
-    4. (d)
-      What are the relations of $P,r,K$ leading to the monotone dependency of the total population on the diffusion coefficient?
-    5. (e)
       Prove or disprove the conjecture that, once $r=f(K/P)$, where $f$ is an increasing function, the total population for large diffusion exceeds the total carrying capacity for the convex $f$ and is less for concave $f$.